OpenStudy (anonymous):

A runner accelerates from rest at a rate of 2 m/s^2. How fast is he running at the end of 5 seconds?

OpenStudy (anonymous):

a = v/t. a= 2m/s^2. t= 5 secounds. solve for v

OpenStudy (anonymous):

i do believe it will be 10m/s at the end of the 5 secounds

OpenStudy (anonymous):

yes, it will be 10 m/s(5*2).

OpenStudy (anonymous):

definately it will be 10m/s